the greater of 2 consecutive even integers is 20 more than twice the smaller. find the integers
Answer by richwmiller(17219)

(n+2)=20+2n
n+2=2n+20
n=2n+18
-n=18
n=-18
n+2=-16
check
-16=20+2(-18)
-16=20-36
ok